Find a quadratic equation for each root.
a)(2/3,-5)
b)(-1/2,-3/4)
c)(0,-3/2)
d)2+root5,2-root5

Answer by edjones(8007)

a)
x=2/3, x=-5
(x- 2/3)(x+5) =0
x^2 + 13x/3 - 10/3 = 0
.
Ed